OpenStudy (gabylovesyou):

If A = (1, 4) B = (3, 1) and C = (6, 3), classify the the following triangle as either equilateral or right. Type the word equilateral or right in the blank. @ganeshie8

OpenStudy (gabylovesyou):

what formula should be used to find this ?

OpenStudy (gabylovesyou):

@dumbcow

OpenStudy (hoblos):

try calculating the length of the 3 sides if they are equal, it would be equilateral if not, you have to check if they apply Pythagoras theorem

OpenStudy (gabylovesyou):

using the slope formula ?

OpenStudy (hoblos):

 Pythagoras theorem : c^2 = a^2 + b^2

OpenStudy (gabylovesyou):

its an equilateral triangle

OpenStudy (hoblos):

all the sides are equal ?

OpenStudy (gabylovesyou):

wait no i think its a right ;/ idk ..

OpenStudy (hoblos):

do you know how to calculate the distance between two points ?

OpenStudy (gabylovesyou):

yes

OpenStudy (hoblos):

so calculate AB=? AC=? BC=?

OpenStudy (gabylovesyou):

AB - 13 BC - 13 CA - 26

OpenStudy (gabylovesyou):

so its a right ?

OpenStudy (hoblos):

they are actually AB = √13 BC = √13 CA = √26 right ?

OpenStudy (gabylovesyou):

yea

OpenStudy (hoblos):

so let's see if it's right is CA^2 = AB^2 + BC^2 ?

OpenStudy (gabylovesyou):

no

OpenStudy (phi):

I would plot it. If the choice is either equilateral or right triangle, you can tell by eyeball

OpenStudy (gabylovesyou):

jk yes haha

OpenStudy (hoblos):

CA^2 = ?

OpenStudy (gabylovesyou):

26 and @phi looks like a right

OpenStudy (phi):

If you want to prove it, find the lengths if each side and show pythagoras works or find the slope of the two lines that form the right angle and see if the slopes are negative reciprocals

OpenStudy (gabylovesyou):

which r the two lines that form the right angle?

OpenStudy (phi):

lines AB and BC

OpenStudy (gabylovesyou):

the slopes dont match -1.5 and 0.66666667

OpenStudy (phi):

-3/2 and ⅔ it's easier to see if you multiply one by -1 and then flip it, you get the other *negative reciprocals*

OpenStudy (phi):

which means the lines are perpendicular (form a 90º angle) so you have proven its a right triangle.